# Reconstructing a Heavy Neutral Lepton at the LHC

### v3

@@ Abstract
[…]
@@ 1 Introduction
Heavy neutral leptons (HNL) provide the simplest UV completion of the Standard Model (SM) that is able to generate the dim-5 Weinberg operator [1], which arguably makes them the most motivated scenario for BSM physics [2, 3, 4, 5, 6, 7, 8]. Let us briefly discuss a minimal setup (see also the majoron model in [9] or the dark matter model in [10]) using 2-component spinors***In this notation $\chi$ is a left-handed spinor and its conjugate contravariant $\bar{\chi}$ is right handed (we omit Lorentz indexes). For example, to define the electron we need two left handed spinors ($e$ and $e^{c}$) of opposite charge; a mass term reads $m\left(ee^{c}+\bar{e}\bar{e^{c}}\right)=m\,\bar{\Psi}_{e}\Psi_{e}$, where $\Psi_{e}$ is the 4-component electron in the chiral representation of $\gamma^{\mu}$: $\Psi_{e}=\begin{pmatrix}e\\ \bar{e^{c}}\end{pmatrix}$ and $\bar{\Psi}_{e}=\left(e^{c}\;\bar{e}\right)$..
@@
The mass $\mu_{1}$ increases in one unit the rank of ${\cal M}$ and defines an inverse seesaw with $m_{\nu_{3}}\approx\mu_{1}(m/M)^{2}$ [11]. The term $\mu_{2}$ does not give mass to a second standard neutrino, it just breaks the degeneracy between the two (now Majorana) heavy neutrinos. The term $\mu_{3}=\tilde{y}_{\nu}v/\sqrt{2}$ is then necessary to give a mass $m_{\nu_{2}}\approx\mu_{3}^{2}/\mu_{1}$ to $\nu^{\prime}_{2}$, with $\nu^{\prime}_{1}$ staying massless. Alternatively, $\mu_{3}$ could be forbidden by a discrete symmetry and $m_{\nu_{2}}$ obtained by adding a second heavy neutrino pair (i.e., an inverse seesaw for the two massive neutrino families).
@@
The possibility of two Majorana HNLs at the TeV scale with sizeable heavy-light mixings (e.g., $m=10$ GeV, $\mu_{1,2}\approx 1$ TeV and $M=0$) requires a fine tuning ($m^{2}/\mu_{2}+\mu^{\prime 2}_{3}/\mu_{1}\approx 10^{-8}m^{2}/\mu_{2}$) that is not stable under radiative corrections. If one of the HNLs is significantly heavier than the other one, the cancelation forces its mixing to be smaller and the model reduces to a single Majorana HNL mixed with a combination of the standard neutrinos. Therefore, although throughout our analysis we will assume a quasi-Dirac HNL $N$, we will also discuss this fine tuned Majorana case.
The heavy-light mixing $s_{\alpha}$ defines then the couplings of $\nu_{\ell}$ and $N$ to the $W^{\pm}$ and $Z$ bosons (we drop the prime to denote mass eigenstates). In particular, $\nu_{\ell}$ will see its gauge couplings reduced by a factor of $c_{\alpha}$ whereas $N$ will now couple to the $W$ with a strength proportional to $V_{\ell N}\approx s_{\alpha}$. The model also implies heavy–light couplings both to the $Z$ boson ($\propto s_{\alpha}c_{\alpha}$) and to the Higgs boson ($y_{\nu}=\sqrt{2}s_{\alpha}M/v$).
Collider bounds on $|V_{\ell N}|^{2}$ may be obtained at energies below or above $m_{N}$. At lower energies the HNL is not produced and the bounds are based on observables like the universality of weak interactions, precision observables related to $\mu$ decays (the mixing with $\nu_{\mu}$ changes the definition of $G_{\mu}$), the invisible width of the $Z$ boson or one-loop flavor changing processes like $\mu\to e\gamma$. A global fit of these observables sets limits on $|V_{\ell N}|^{2}$ ranging from $10^{-2}$ to $10^{-4}$, being the most stringent ones for $\ell=\mu$ [12, 13, 14].
@@
Here we will discuss higher energy processes with the direct production of the HNL. The current collider limits for $m_{N}>M_{W}$ have been obtained by CMS in the trilepton channel at the LHC [15, 16] and by L3 at LEP [17] (see the dominant diagrams in Fig. 1), in this second case only for $\ell=e$. We will focus on the trilepton channel. We will show that, although the longitudinal momentum of the $\nu$ escapes detection, the mass $m_{N}$ of the HNL can be reconstructed. We will also comment on the possible effects introduced by a HNL in Higgs searches at the LHC. In particular, notice (see Fig. 2) that the dilepton plus missing $p_{\rm T}$ signal from $h\to WW^{*}$ observed at CMS coincides with the one from $h,Z\to\nu N$, although the kinematics in each process is obviously different.
@@ 2 HNL at the LHC
To study the signal of a HNL at the LHC we first produce an UFO file [18] with FeynRules [19] that includes the new interactions, then we generate signal and background events with MadGraph5 [20] + Pythia [21], and finally we analyze the results with routines based on FasJet [22], ROOT [23] and HepMC [25].
Let us consider a $90$, $110$ or $130$ GeV HNL with, in all cases, a mixing $|V_{\mu N}|^{2}=10^{-3}$ with the muon flavorAs we will see, for this mixing and the current luminosity the LHC can not access larger masses.. For 138 fb-1 at 13 TeV, our simulation estimates, respectively, 1021, 317 or 153 HNL events giving two opposite sign muons, an electron and $\vec{p}_{\rm T}^{\rm\,miss}$:
@@
In the simulation we have included the events where the final electron comes from the leptonic decay of a tau lepton (i.e., the $\ell^{\prime+}$ in Fig. 1 may be an $e^{+}$ or a $\tau^{+}$ decaying into $\bar{\nu}_{\tau}\,e^{+}\nu_{e}$). We then apply the cuts defined in [16], namely, the three leptons must be isolated and with a $p_{\rm T}>10$ GeV (highest $p_{\rm T}$ above 55 GeV), a pseudorapidity $|\eta|<2.4\,(2.5)$ for the muons (electron), and an invariant mass $m_{\mu\mu}>5$ GeV with $|m_{\mu\mu}-m_{Z}|>15$ GeV and $|m_{\mu\mu e}-m_{Z}|>15$ GeV. We find that 20 events (just 2% of the initial HNL sample) pass the cuts for $m_{N}=90$ GeV, 17 events for $m_{N}=110$ GeV and 14 for $m_{N}=130$ GeV. The low efficiency of the cuts suggests that the CMS analysis has been optimized for larger HNL massesFor $m_{N}=90$ GeV, in particular, we find that a reduction from 10 to 5 GeV in the minimum $p_{T}$ of the least energetic lepton increases the signal by a factor of 1.42 and the background by a factor of 1.12)..
The background, in turn, is dominated by $WZ$ and $ZZ$ production: we obtain that the cuts are satisfied by 285 events, which define around 50% of the total background yield [16] (e.g., these 285 events do not include triboson events, leptons from heavy hadron decays nor photons and jets misidentified as leptons).
The momenta of the two muons and the electron in the HNL events can be measured, whereas the transverse momentum $\vec{p}^{\,\nu}_{\rm T}$ of the neutrino can be obtained from $\vec{p}_{\rm T}^{\rm\,miss}$. In addition, we notice that for $m_{N}>m_{W}$ the $W$ boson decaying into electron plus neutrino is on shell. This can be used to deduce the longitudinal momentum $p_{\rm L}^{\nu}$ of the neutrino (see [26], for example, for an analogous calculation in a different context): it is solution to the quadratic equation
@@
Only one of the two solutions for $p_{L}^{\nu}$ in Eq (5) provides the right reconstruction of $m_{N}$, and we find the lower one to be the best choice. In particular, when the two solutions are significantly different ($m_{2}>1.1\,m_{1}$) this is the right choice 96% of the times, whereas for similar solutions the prescription is successful 84% of the times.
Figure 3: Mass reconstruction in $pp\to\mu^{\pm}\mu^{\mp}e^{\pm}\nu$ (left) and $pp\to\mu^{\pm}\mu^{\mp}\mu^{\pm}\nu$ (right) at 13 TeV, 138 fb-1 for $m_{N}=(90,110,130)$ GeV with $|V_{\mu N}|^{2}=10^{-3}$ and diboson background.
In Fig. 3 we plot the distribution of the reconstructed $m_{N}$ for the 20, 17 and 14 signal events that correspond to the three HNL masses together with the 285 background events (these distributions have been obtained with a normalized sample 20 times larger). Although our estimate does not include a simulation of the detector response, the peak obtained for a mixing $|V_{\mu N}|^{2}=10^{-3}$ could be significant at the lower values of $m_{N}$.
@@
In this case, the muon needed to reconstruct $m_{N}$ appears together with a second muon with the same charge. We notice, however, that the production of a HNL (see Fig. 1) is favored by the collision of a valence quark and a sea antiquark in the initial protons. This tends to give a larger $|p_{\rm L}|$ to the muon in the primary vertex and, together with the cut $p^{\rm miss}_{\rm T}>55$ GeV, implies that $m_{N}$ must be reconstructed with the muon of lower energy. The Monte Carlo simulation shows that this prescription gives the right choice 78% of the times for $m_{N}=90$ GeV or 87% for $m_{N}=130$ GeV.
Figure 4: Mass reconstruction in $pp\to e^{\pm}e^{\mp}\mu^{\pm}\nu$ (left) and $pp\to e^{\pm}e^{\mp}e^{\pm}\nu$ (right) at 13 TeV, 138 fb-1 for $m_{N}=(90,110,130)$ GeV with $|V_{eN}|^{2}=10^{-3}$ and diboson background.
Therefore we consider as well this trimuon signal. Again, we find that only 2% of the HNL events pass the cuts for $m_{N}=90$ GeV or 6% at $m_{N}=130$ GeV. The reconstruction of $m_{N}$ is done with the criteria described above but changing the electron by the less energetic muon among the two with same charge. The event distribution for this observable is shown in Fig. 3-right. Combined with the dimuon plus electron events and normalizing the background by a factor of 2§§§As shown in [16], triboson/top quark events, nonprompt leptons and missidentified jets/photons account approximately for $50\%$ of the background., we estimate that at the current luminosity the bounds on $|V_{\mu N}|^{2}$ at CMS and ATLAS could be around $4.0\times 10^{-4}$ for $m_{N}=90$ GeV or $8.4\times 10^{-4}$ for $m_{N}=130$ GeV. These limits are set at 95% confidence level using the $CL_{s}$ prescription [24]. An analogous analysis for $pp\to e^{\pm}N\to e^{\pm}e^{\mp}\ell^{\prime\pm}\nu$ (in Fig. 4) suggests very similar bounds for the mixing with the electron flavor: $|V_{eN}|^{2}<3.8\times 10^{-4}$ or $8.0\times 10^{-4}$ for a HNL mass of $90$ or $130$ GeV, respectively.
@@ 3 Majoranas and other production channels
We can also reconstruct $m_{N}$ for a heavy Majorana singlet. This case implies $L$-violating processes with $p_{\rm T}^{\rm miss}$, two muons (or electrons) with the same charge plus an electron (or muon) of opposite charge (see Fig. 1), e.g.,
@@
To reconstruct $m_{N}$ we need to choose among the two same sign muons. As explained before, the muon from $W^{+}\to\mu^{+}N$ tends to be the most energetic one. Here, in addition, the chirality flip in $N$ favors that the $\mu^{+}$ from $N\to\mu^{+}W^{-}$ is emitted backwards [27], which also reduces its energy: 89% of the times the reconstruction with the less energetic muon provides the right value for $m_{N}=110$ GeV. Again, the efficiency of the cuts seems low: 19 out of 511 HNL events pass the cuts for $m_{N}=90$ GeV or 10 out of 83 for $m_{N}=130$ GeV. We find, however, that the reconstruction of $m_{N}$ works and, together with a total background 50% smaller than in the Dirac case, our results suggest similar limits. For example, at $m_{N}=90$ GeV we estimate
$|V_{eN}|^{2}<4.0\times 10^{-4}\,;\hskip 14.22636pt|V_{\mu N}|^{2}<4.2\times 10^{-4}\,.$ (9)
Finally, we would like to comment on the reconstruction of the HNL mass also in other channels. In particular, Higgs searches $h\to WW^{*}\to\mu^{+}\nu\,e^{-}\bar{\nu}$ (dilepton of different flavor plus missing $p_{\rm T}$) [28, 29, 30, 31, 32] with slightly different cuts could be sensitive to the presence of the HNL: $Z(h)\to\nu N\to\nu\,\mu^{+}W^{-}\to\nu\,\ell^{+}\bar{\nu}\,e^{-}$.
Take $m_{N}=85$ GeV (the argument applies to any masses between $m_{W}$ and around $100$ GeV) and $|V_{\mu N}|^{2}=10^{-3}$. Before the cuts described in [33], we estimate 9,950 HNL events versus 11,320 Higgs events. We notice that the low value of $m_{N}$ strongly favors that the s-channel $Z$ boson in Fig. 2 is near the mass shell, implying that the neutrino from $Z\to\nu N$ will carry little $p_{\rm T}$. We can then assume that most of the missing $p_{\rm T}$ is carried by the second neutrino and, as before, use that the parent $W$ is on shell to find $p^{\nu}_{L}$ and reconstruct $m_{N}$. We find, however, that with the cuts optimized for Higgs searches very few $N$ events are selected: the ones used in [33] keep only 0.7% of the 9,950 events, versus 11% of the 11,320 Higgs events. The basic reason is that the charged lepton from $N\to\mu^{+}W^{-}$ tends to have a $p_{\rm T}$ below the 15 GeV required to the subleading lepton. If we relax this minimum $p_{\rm T}$ to 5 GeV and impose that the electron is more energetic than the muon we would keep 1.1% of the $N$ sample (106 events), while the fraction of $h\to WW^{*}$ events passing the cuts would be reduced to 4.4% (500 Higgs events). These frequencies suggest a possible complementarity of Higgs physics and $N$ searches in some mass region.

### v4

@@ Abstract
Heavy lepton singlets $N$ slightly mixed with a standard neutrino $\nu_{\ell}$ are usually searched for at the LHC in the trilepton plus $p_{\rm T}^{\rm miss}$ channel: $pp\to W^{+}\to\ell^{+}N$ with $N\to\ell^{-}W^{+}\to\ell^{-}\ell^{\prime+}\nu$. We show that, although the longitudinal momentum of the final $\nu$ escapes detection, the mass of the heavy lepton can be reconstructed. While this possibility has not been considered in recent LHC searches, we find that the search for a mass peak could systematically improve the current collider bounds on the mixing $|V_{\ell N}|^{2}$ for any mass $m_{N}\geq M_{W}$.
@@ 1 Introduction
Heavy neutral leptons (HNL) provide the simplest UV completion of the Standard Model (SM) that is able to generate the dim-5 Weinberg operator [1], which arguably makes them the most motivated scenario for BSM physics [2, 3, 4, 5, 6, 7, 8]. Let us briefly discuss a minimal setup (see also the models in [9, 10]) using 2-component spinors.
@@
The mass $\mu_{1}$ increases in one unit the rank of ${\cal M}$ and defines an inverse seesaw with $m_{\nu_{3}}\approx\mu_{1}(m/M)^{2}$ [11]. The term $\mu_{2}$ does not give mass to a second standard neutrino, it just breaks the degeneracy between the two (now Majorana) heavy neutrinos. The term $\mu_{3}=\tilde{y}_{\nu}v/\sqrt{2}$ is then necessary to give a mass $m_{\nu_{2}}\approx\mu_{3}^{2}/\mu_{1}$ to $\nu^{\prime}_{2}$, with $\nu^{\prime}_{1}$ staying massless. Alternatively, $\mu_{3}$ could be forbidden by a discrete symmetry and $m_{\nu_{2}}$ obtained by adding a second heavy neutrino pair (i.e., an inverse seesaw for the two massive neutrino families, with the third neutrino massless).
@@
The possibility of two Majorana HNLs at the TeV scale with sizeable heavy-light mixings (e.g., $m=10$ GeV, $\mu_{1,2}\approx 1$ TeV and $M=0$) requires a fine tuning ($m^{2}/\mu_{2}+\mu^{\prime 2}_{3}/\mu_{1}\approx 10^{-8}m^{2}/\mu_{2}$) that is not stable under radiative corrections. If one of the HNLs is significantly heavier than the other one, the cancelation forces its mixing to be smaller and the model reduces to a single Majorana HNL mixed with a combination of the standard neutrinos. Therefore, although throughout our analysis we will assume a quasi-Dirac HNL $N$, we will also comment on this fine tuned Majorana case.
The heavy-light mixing $s_{\alpha}$ defines then the couplings of $\nu_{\ell}$ and $N$ to the $W^{\pm}$ and $Z$ bosons (we drop the prime to denote mass eigenstates). In particular, $\nu_{\ell}$ will see its gauge couplings reduced by a factor of $c_{\alpha}$ whereas $N$ will now couple to the $W$ with a strength proportional to $V_{\ell N}\approx s_{\alpha}$. The model also implies heavy–light couplings both to the $Z$ boson ($\propto s_{\alpha}c_{\alpha}$) and to the Higgs boson ($y_{\nu}=\sqrt{2}s_{\alpha}M/v$). We will consider the case where $N$ couples to a single generation of SM neutrinos [12, 8].
Collider bounds on $|V_{\ell N}|^{2}$ may be obtained at energies below or above $m_{N}$. At lower energies the HNL is not produced and the bounds are based on observables like the universality of weak interactions, precision observables related to $\mu$ decays (the mixing with $\nu_{\mu}$ changes the definition of $G_{F}$), the invisible width of the $Z$ boson or one-loop flavor changing processes like $\mu\to e\gamma$. A global fit of these observables sets PMNS unitarity limits on $|V_{\ell N}|^{2}$ ranging from $10^{-2}$ to $10^{-4}$, being the most stringent ones for $\ell=\mu$ [13, 14, 15, 16].
@@
Here we will discuss higher energy processes with the direct production of the HNL. The signatures of such models at colliders have been studied extensively [17, 18, 19, 20, 21, 22, 23, 24, 25]. The current bounds on $|V_{\ell N}|^{2}$ for $m_{N}>M_{W}$ have been obtained by CMS in the trilepton channel at the LHC [26, 27] and by L3 at LEP [28], in this second case only for $\ell=e$. We will focus on the trilepton channel, a process dominated by the charged-current Drell–Yan $q\bar{q}^{\prime}\to W^{\pm}\to Nl^{\pm}$ [29, 30] (see Fig. 1). In our analysis we will also include vector boson fusion $q\gamma\to Nl^{\pm}q^{\prime}$ [31, 32, 33], a process giving a significant contributions for masses $m_{N}\geq 500$ GeV. We will show that, although the longitudinal momentum of the $\nu$ escapes detection, the mass $m_{N}$ of the HNL can be reconstructed.
We will comment as well on the possible effects introduced by an HNL in Higgs searches at the LHC. In particular, notice (see Fig. 2) that the dilepton plus missing $p_{\rm T}$ signal from $h\to WW^{*}$ observed at CMS coincides with the one from $h,Z\to\nu N$, although the kinematics in each process is obviously different.
@@ 2 HNL at the LHC
To estimate the possible signal of an HNL at the LHC, we first produce a next-to-leading-order (NLO) UFO file [34] with the new vertices using the Mathematica packages FeynRules [35], NLOCT [36] and FeynArts [37]. Then, we generate signal and background events at NLO using MadGraph5_aMC@NLO v2.9.7 [38] interfaced with Pythia v8.310 [39], MadSpin [40] and Delphes v3.2.0 [41]. In Delphes we have used the CMS simulation card, including a photon conversion module based on the characteristics of the detector [42, 43, 44, 45]. For the parton distribution functions (PDFs), we use NNPDF3.1 [46] in all the cases except for vector boson fusion signal process, where we use NNPDF31.luxQED [47, 48, 49] (both as implemented in lhapdf [50]). For the merging between jets from matrix elements and from parton showers, we have used the FxFx matching scheme [51]. Finally, we have analyzed the results with custom routines based on FasJet [52], ROOT [53] and HepMC [54].
After we generate the events, we apply the cuts defined in [27] for the high-mass regime, with at least one opposite-sign same-flavour (OSSF) lepton pair and
  1. 1.
    The three leptons must be isolated as defined in [55], with $p_{\rm T}(\ell_{1})>55$ GeV, $p_{\rm T}(\ell_{2})>15$ GeV, $p_{\rm T}(\ell_{3})>10$ GeV.
  2. 2.
    No $b$ jets.
  3. 3.
    Any OSSF pair must satisfy $m(\ell^{+}\ell^{-})>5$ GeV and $\left|m(\ell^{+}\ell^{-})-m_{Z}\right|>15$ GeV.
  4. 4.
    The trilepton invariant mass is $\left|m(3\ell)-m_{Z}\right|>15$ GeV.
To illustrate our results, let us consider three benchmarks masses $m_{N}$ in the 90909090170170170170 GeV range and two more in the higher mass region 400400400400600600600600 GeV. In all the cases we assume a mixing $|V_{\mu N}|^{2}=10^{-2}$ with the muon flavor (the cross sections are proportional to $|V_{\mu N}|^{2}$) and we take a luminosity of 138 fb-1 at 13 TeV. We first focus on the channel
@@
[…]
The background, in turn, is made up of $WZ$ (37%), $ZZ$ (1%), $Z\gamma$ (38%), nonprompt ($Z$+jets, $t\bar{t}$+jets, 20%) and other ($WWW$, $WWZ$, $ZZ$, $t\bar{t}W$, $t\bar{t}Z$, $t\bar{t}\gamma$, 4%). We estimate 54840 events that are reduced to 1259 background events after the cuts. These frequencies are consistent (within a 20%) with the ones used in [27].
$m_{N}$ [GeV] 90 130 170 400 600
$\mu^{\pm}\mu^{\mp}e^{\pm}$ events 858 297 101 4.01 1.78
events after cuts 148 119 72.5 3.87 1.73
(17.2%) (40.0%) (71.8%) (96.5%) (97.2%)
Table 1: Signal events in the channel $\mu^{\pm}\mu^{\mp}e^{\pm}\nu$ before and after cuts (138 fb-1 at 13 TeV) for several HNL masses and $|V_{\mu N}|^{2}=10^{-2}$.
Now, the detector provides the momenta of the two muons and of the electron in all the signal events, whereas the transverse momentum $\vec{p}^{\,\nu}_{\rm T}$ of the neutrino can be obtained from $\vec{p}_{\rm T}^{\rm\,miss}$. In addition, we notice that for $m_{N}>m_{W}$ the $W$ boson decaying into electron plus neutrino is on shell. This can be used to deduce the longitudinal momentum $p_{\rm L}^{\nu}$ of the neutrino (see [56] for an analogous calculation in a different context): it is solution to the quadratic equation
@@
We find, however, that after the detector simulation 8% of the signal events for $m_{N}=90$ GeV or 13% for 600 GeV have no real solution to Eq (5). This introduces an effective extra cut in the signal that, interestingly, is stronger on the background: 25% of the background events are cut by the requirement of having a solution to Eq (5). For the rest of events the equation implies a two-fold degeneracy, and we find the lower solution to be the best choice 87% of the times.
In Fig. 3 we plot the reconstruction of different values of $m_{N}$ in the low and high mass regions together with the background.
Figure 3: Reconstruction of $m_{N}$ from $pp\to\mu^{\pm}\mu^{\mp}e^{\pm}\nu$ in the low (left) and high (right) mass regions for $|V_{\mu N}|^{2}=0.01$ and $0.1$, respectively, together with the background (from top to bottom, the contributions from $Z\gamma$, $WZ$, $ZZ$ and nonprompt+other separated by dashes).
@@
We plot the number of events before and after cuts in Table 2. This trimuon channel has an estimated background of 31321 events (60% from $WZ$, 36% from $ZZ$, 3% nonprompt and below1% from other) that is reduced to 642 events after cuts.
In this case, the muon needed to reconstruct $m_{N}$ appears together with a second muon with the same charge. We notice, however, that the production of an HNL (see Fig. 1) is favored by the collision of a valence quark and a sea antiquark in the initial protons. This tends to give a larger $p_{\rm T}$ to the muon in the primary vertex and implies that $m_{N}$ must be reconstructed with the muon of lower $p_{\rm T}$. We find that this prescription works 92% of the times for $m_{N}=130$ GeV or 77% for $m_{N}=400$ GeV.
Therefore we consider as well this trimuon channel. The reconstruction is done with the same criteria as before but changing the electron by the muon with less $p_{\rm T}$ among the two with same charge. The distribution for the reconstructed $m_{N}$ is shown in Fig. 3 in the two mass regimes. Combining the $\mu\mu e$ and the $\mu\mu\mu$ channels, we obtain an estimate of the expected LHC limits at 95% confidence level (we use the $CL_{s}$ prescription [57]). We find bounds (see Table 3) that go from $|V_{\mu N}|^{2}<7.8\times 10^{-4}$ for $m_{N}=90$ GeV to $|V_{\mu N}|^{2}<6.2\times 10^{-2}$ for $m_{N}=600$ GeV.
$m_{N}$ [GeV] 90 130 170 400 600
$\mu^{\pm}\mu^{\mp}\mu^{\pm}$ events 981 336 112 4.39 1.90
events after cuts 164 92.3 53.7 4.13 1.85
(16.7%) (27.4%) (47.9%) (94.1%) (97.3%)
Table 2: Signal events in the channel $\mu^{\pm}\mu^{\mp}\mu^{\pm}\nu$ before and after cuts (138 fb-1 at 13 TeV) for several HNL masses and $|V_{\mu N}|^{2}=10^{-2}$.
We can estimate the possible relevance of this observable in current collider searches. In a recent analysis [27], CMS set bounds using boosted decision trees (BDT) trained with several kinematic observables (mainly $\Delta\text{R}\left[\text{min }m(l^{+}l^{-})\right]$, $m_{T}$, $p_{T}(l_{3})$ and $p_{T}^{\text{miss}}$). Our results indicate that this optimized combination of observables provides a less efficient signal-background discrimination than $m_{N}$, especially in the low mass region: $|V_{\mu N}|^{2}<2.2\times 10^{-3}$ (versus $7.8\times 10^{-4}$) for $m_{N}=90$ GeV or $|V_{\mu N}|^{2}<5.6\times 10^{-3}$ (versus $2.6\times 10^{-4}$) for $m_{N}=130$ GeV. Therefore, significantly better bounds could be expected if $m_{N}$ were added to the pool of observables training the BDT.
Figure 4: Reconstruction of $m_{N}$ from $pp\to\mu^{\pm}\mu^{\mp}\mu^{\pm}\nu$ in the low (left) and high (right) mass regimes for $|V_{\mu N}|^{2}=0.01$ and $0.1$, respectively, together with the background (from top to bottom, the contributions from $WZ$, $ZZ$ and nonprompt+other separated by dashes).
$m_{N}$ [GeV] 90 130 170 400 600
$|V_{\mu N}|^{2}$ limit $7.8\times 10^{-4}$ $2.6\times 10^{-3}$ $4.9\times 10^{-3}$ $3.8\times 10^{-2}$ $6.2\times 10^{-2}$
Table 3: Expected bounds on $|V_{\mu N}|^{2}$ by combining the channels $\mu^{\pm}\mu^{\mp}e^{\pm}\nu$ and $\mu^{\pm}\mu^{\mp}\mu^{\pm}\nu$ for 138 fb-1 at 13 TeV.
An analogous analysis can be done for the mixing $|V_{eN}|$,
$pp\to W^{\pm}\to e^{\pm}N\to e^{\pm}e^{\mp}\,W^{\pm}\to e^{\pm}e^{\pm}\ell^{\mp}\,\nu\,.$ (8)
Our results suggests similar bounds in this case. For example, we obtain a 20% smaller signal after cuts at $m_{N}=130$ GeV within a 10% smaller background, implying an expected bound around $|V_{eN}|^{2}<2.9\times 10^{-3}$ for this HNL mass.
@@ 3 Majoranas and other production channels
We can also reconstruct $m_{N}$ for a heavy Majorana singlet. We focus in the $L$-violating processes with $p_{\rm T}^{\rm miss}$, two muons (or electrons) with the same charge plus an electron (or muon) of opposite charge (see Fig. 1), e.g.,
@@
To reconstruct $m_{N}$ we need to choose among the two same sign muons. As explained before, the muon from $W^{+}\to\mu^{+}N$ tends to be the one with larger $p_{\rm T}$ Here, in addition, the chirality flip in $N$ favors that the $\mu^{+}$ from $N\to\mu^{+}W^{-}$ is emitted backwards [58], which also reduces its energy: for $m_{N}=130$ GeV, 89% of the times the reconstruction with the muon that has lower $p_{\rm T}$ provides the right value.
Although we find that the reconstruction of $m_{N}$ works, the number of signal events in the channel $\mu^{+}\mu^{+}e^{-}$ is half the one in $\mu^{+}\mu^{-}e^{+}$ for the Dirac case, plus here the trimuon channel does not provide an efficient reconstruction of $m_{N}$. This, together with a background that is now only composed by $t\bar{t}$+jets, $WWW$ and $t\bar{t}W$, imply similar limits: $|V_{\mu N}|^{2}<2.1\times 10^{-3}$ and $|V_{eN}|^{2}<2.5\times 10^{-3}$ for $m_{N}=130$ GeV.
Finally, we would like to comment on the reconstruction of the HNL mass also in other channels. In particular, we would like to argue that Higgs searches $h\to WW^{*}\to\mu^{+}\nu\,e^{-}\bar{\nu}$ (dilepton of different flavor plus missing $p_{\rm T}$) [59, 60, 61, 62, 63] with slightly different cuts could be sensitive to the presence of the HNL:
$pp\to Z(h)\to\nu N\to\nu\,\mu^{+}W^{-}\to\nu\,\ell^{+}\bar{\nu}\,e^{-}.$ (10)
Take $m_{N}=85$ GeV (the argument applies to any masses between $m_{W}$ and around $100$ GeV) and $|V_{\mu N}|^{2}=10^{-3}$. Before the cuts described in [64], we estimate 9950 HNL events versus 11320 Higgs events. We notice that the low value of $m_{N}$ strongly favors that the s-channel $Z$ boson in Fig. 2 is near the mass shell, implying that the neutrino from $Z\to\nu N$ will carry little $p_{\rm T}$. We can then assume that most of the missing $p_{\rm T}$ is carried by the second neutrino and, as before, use that the parent $W$ is on shell to find $p^{\nu}_{L}$ and reconstruct $m_{N}$. We find, however, that with the cuts optimized for Higgs searches very few $N$ events are selected: the cuts used in [64] keep only 0.7% of the 9950 events, versus 11% of the 11320 Higgs events. The basic reason is that the charged lepton from $N\to\mu^{+}W^{-}$ tends to have a $p_{\rm T}$ below the 15 GeV required to the subleading lepton. If we relax this minimum $p_{\rm T}$ to 5 GeV and impose that the electron is more energetic than the muon we would keep 1.1% of the initial $N$ sample (106 events), while the fraction of $h\to WW^{*}$ events passing the cuts would be reduced to 4.4% (500 Higgs events). These estimates suggest a possible complementarity of Higgs physics and $N$ searches in this mass region.